OpenStudy (spring98):

The incorrect work of a student to solve an equation 2(y + 8) = 4y is shown below: Step 1: 2(y + 8) = 4y Step 2: 2y + 10 = 4y Step 3: 2y = 10 Step 4: y = 5 Which of the following explains how to correct Step 2 and shows the correct value of y? (5 points) The equation should be y + 8 = 4y after division by 2; y = 4 The equation should be y + 8 = 4y after division by 2; y = 6 2 should be distributed as 2y + 16; y = 4 2 should be distributed as 2y + 16; y = 8

OpenStudy (spring98):

Can you help me on this one"?

OpenStudy (haseeb96):

yeah figure out the mistake in 2nd step

OpenStudy (spring98):

i tried but i couldn't figure out the right answer

OpenStudy (haseeb96):

2(y+8)=4y 2y+16=4y 4y-2y=16 2y=16 y=16/2 y=8 this is your correct answer

OpenStudy (haseeb96):

now compare it with the above explanation

OpenStudy (spring98):

so that means that the last one is the answer?

OpenStudy (haseeb96):

absolutely

OpenStudy (spring98):

thank you very much! some other guy could not explain it to me as good as you did. :)

OpenStudy (haseeb96):

I am just a little boy. GOD has born me to help the people .